OpenStudy (anonymous):

simplify: ln(e^3/x^4)

OpenStudy (anonymous):

i've gotten 3/x^4 - is that correct?

OpenStudy (anonymous):

what happened to ln?

OpenStudy (anonymous):

remember, \[\ln a/b = \ln a - \ln b \]\[lna^b = b lna\]

OpenStudy (anonymous):

3 - 4ln(x)? or does lnx^4 become -x^4?

OpenStudy (anonymous):

\[3-4\ln(x)\]) is correct

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem.